I don't understand the following math problem and would appreciate any help.
7/8 = n+2/3

Can someone please explain?

Thank you!!

Answer by rfer(16322)

use common denominators
7/8=n+2/3
subtract 2/3 from both sides
7/8-2/3=n
common denominators
21/24-16/24=n
subtract and reduce if you can, can't here though
5/24=n